OpenStudy (anonymous):

MEDAL!!!!! If it takes Wilbur 3 hours to rake the leaves in a neighbors’ backyard, and it takes his friend, Orville 6 hours to rake the same amount of leaves, how long would it take Wilbur and Orville to rake the leaves in a neighbors’ backyard together?

OpenStudy (anonymous):

The equation is 1 over 3 + 1 over 6 = 1 over x I multiplied each part by 6x and wound up getting 2x + x = 6. How do I solve from here?

OpenStudy (anonymous):

So you have 1/3 + 1/6 = 1/x

OpenStudy (anonymous):

Would the common multiple be 6x?

OpenStudy (anonymous):

Next RULE: combine all numbers on one side

OpenStudy (anonymous):

and the variable on the other side

OpenStudy (anonymous):

So first, leave the variable alone and combine "1/3 + 1/6" on the left side

OpenStudy (anonymous):

How do I do that?

OpenStudy (anonymous):

"1/3 + 1/6" = "2/6 + 1/6"

OpenStudy (anonymous):

=3/6

OpenStudy (anonymous):

=1/2

OpenStudy (anonymous):

Now we have 1/2 = 1/x

OpenStudy (anonymous):

Next RULE: solve for x

OpenStudy (anonymous):

To do so, we would normally need to isolate x by multiplying or dividing. Here, we know x by just looking at the equation above

OpenStudy (anonymous):

Solve from the 1/2, you mean?

OpenStudy (anonymous):

Hint: 1/___ =1/____ 1/2=1/2 1/3=1/3 1/4=1/4 1/5=1/5

OpenStudy (anonymous):

1/2 = 1/x 1/2 = 1/2 x=2

OpenStudy (anonymous):

Oh, I see, because then 2 makes each of the fractions equal zero?

OpenStudy (anonymous):

Lets simplify this: The RULE we use to solve this: to isolate a variable, subtract, add, multiply and divide to get numbers on one side and variables on the other side

OpenStudy (anonymous):

Lets apply this rule

OpenStudy (anonymous):

So you have 1/3 + 1/6 = 1/x

OpenStudy (anonymous):

Yes, okay..

OpenStudy (anonymous):

combine "1/3 + 1/6" on the left side

OpenStudy (anonymous):

2/6 + 1/6 on the left side

OpenStudy (anonymous):

=3/6 on the left side =1/2 on the left side

OpenStudy (anonymous):

We have 1/2 = 1/x. Now, lets isolate the variable x

OpenStudy (anonymous):

to do so, multiply each side by x

OpenStudy (anonymous):

giving us x/2 = 1

OpenStudy (anonymous):

Ok, so in order for the two sides to equal the same, the right has to be a 1/2 too, no?

OpenStudy (anonymous):

next, multiply each side by 2

OpenStudy (anonymous):

that is one way to look at it-but it is a shortcut so I am re-doing the problem...disregard my prior explanation

OpenStudy (anonymous):

so multiplying each side by 2, we get x=2

OpenStudy (anonymous):

Just use the rule "to isolate a variable, subtract, add, multiply and divide to get numbers on one side and variables on the other side"

OpenStudy (anonymous):

Ohhhh, okay, I understand it better now; thanks:)

OpenStudy (anonymous):

One question though, how did the 1/3 turn into 2/6 in the beginning?

OpenStudy (anonymous):

multiply 1/3 x 2/2

OpenStudy (anonymous):

An in-person drawing would work better here...I will draw the explanation later if you do not understand (gotta go now unfortunately)

OpenStudy (anonymous):

to make the LCD for the left side 6

OpenStudy (anonymous):

That's okay, but why was 2/2 multiplied by the 1/3 and not the 1/6?